I have asked for help on these and have not gotten any response.Can anyone help? Give the domain and range of the following. A. (2,5);(9,8) B. y=3x C.y=[2x] D.x=2+3y I am working out of home school books they do not give a good explanation . I have tried books from the library and cant find info on this part of algebra. Please help if you can . I am not asking anyone to do my work for me. I really need help.
Answer by edjones(8007)

Domain relates to x; range to y.
A. There is a straight line between the 2 points given. The domain is from 2 to 9 inclusive or[2,9], the range is [5,9]
B. The domain is all real numbers. The range is from -infinity to infinity.
C. y=|2x| The domain is all real numbers. The range is from 0 to infinity or [0,infinity) because 2x can't be negative.
D. x=2+3y
3y=x-2
y=(x-2)/3
Domain is all real numbers. Range is (-infinity,infinity).
.
Hope that helps.
.
Ed